OpenStudy (aaronandyson):

on what sum will the compound interest for 2 years at 5% p.a be $205

OpenStudy (aaronandyson):

@imqwerty

OpenStudy (welshfella):

Let the original sum be S, then the sum after 2 years will be S + 205 so;- S + 205 = S(1 + 0.05)^2 solve for S

OpenStudy (aaronandyson):

@Michele_Laino

Parth (parthkohli):

welsh is right

rvc (rvc):

yep

OpenStudy (aaronandyson):

S + 205 = S(1 + 0.05)^2 S + 205 = 1.1025S

OpenStudy (tkhunny):

You're missing an "S" on the RHS.

rvc (rvc):

he has written S on rhs at the end

OpenStudy (tkhunny):

Really?! Wow, I think I need a better screen. If you say so. Okay, then. What's preventing you from finishing the solution?

OpenStudy (aaronandyson):

Im getting 0.000041 as the answer

OpenStudy (aaronandyson):

@dan815

OpenStudy (dan815):

S + 205 = S(1 + 0.05)^2

OpenStudy (dan815):